OpenStudy (anonymous):

integral of sqrt(4x-x^2)dx

OpenStudy (anonymous):

Complete the square: \[4x-x^2=-\left(x^2-4x\right)=-\left(x^2-4x+4-4\right)=4-\left(x-2\right)^2\] \[\int\sqrt{4-(x-2)^2}~dx\] You've learned trig subs, correct?

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @SithsAndGiggles Let \(x-2=2\sin u\), so that \(dx=2\cos u~du\): \[\int\sqrt{4-\left(2\sin u\right)^2}~(2\cos u~du)\\ \int\sqrt4\sqrt{1-\sin^2 u}~(2\cos u~du)\\ 4\int\sqrt{1-\sin^2 u}~\cos u~du\] From here, you can simplify. \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

\[\int\limits\limits(\sqrt{4-(x-2)^2} )dx \] \[(x-2)/2*\sqrt{4-(x-2)^2} + 2\sin^{-1} ((x-2)/4) +C\]

OpenStudy (anonymous):

Standard Formula